# On the torsion growth in quadratic number fields for elliptic curves defined over the rationals

### v1

@@ 3. The prime $\ell=2$: The mixed case
Let us continue with the case where a new torsion point $P\in E(K)$ appears, such that a nontrivial multiple $mP$ belongs to $E(\mathbb{Q})$. As mentioned in the Introduction, from [2, Thm. 2], we can reduce ourselves to the following hypothesis:
There exists PE(K)[N]E() such that [2]PE() and $\text{There exists $P\in E(K)[N]\setminus E(\mathbb{Q})$ such that $[2]P\in E(\mathbb{Q})$ and $o(P)=N$},$
@@
Let $E/\mathbb{Q}$ be an elliptic curve and $K=\mathbb{Q}(\sqrt{d})$ a quadratic extension. Let us suppose that there exists $P\in E(K)[4]\setminus E(\mathbb{Q})$ such that $\mathcal{O}\neq[2]P\in E(\mathbb{Q})$. Then if $2|d$, $E$ must have bad reduction at $2$.
@@
where formulas for $b_{i}$ can be found in [11, Ch. III]. It is easy to see that the discriminant of this equation is still $\Delta$. In order to get an equation with integer coefficients, we make the change
@@
By hypothesis, $\mathcal{O}\neq[2]P\in E_{\mathrm{tors}}(\mathbb{Q})$. So there exists $[2]P=Q=(\gamma,0)$ with $\gamma\in\mathbb{Z}$ by Nagell-Lutz theorem. Now, we apply the change of variables
@@
Case I: $E_{\mathrm{tors}}(\mathbb{Q})[2]=\mathcal{C}_{2}$. We can use the following lemma proven in [3, Lemma 13].
@@
which implies that $v_{2}(s)\leq 2$. Now we get a contradiction in each possible case:
I.A: $v_{2}(s)=2$. As $\mathbb{Q}(\sqrt{d})=\mathbb{Q}(\sqrt{A\pm 2s})$ and $2|d$, we have
$2|(A\pm 2s),$
which implies $v_{2}(A-2s)+v_{2}(A+2s)\neq 0$, a contradiction.
I.B: $v_{2}(s)=1$. Equation (6) becomes
(7) $4=v_{2}(A-2s)+v_{2}(A+2s).$
If $v_{2}(A)=2$, we have that $A=4M$, with $M\in\mathbb{Z}$ an odd integer. On the other hand, $2s=4r$ with $r$ an odd integer. Thus $A\pm 2s=4M\pm 4r=4(M\pm r)$, so that $v_{2}(A\pm 2s)\geq 3$. We have $4=v_{2}(A-2s)+v_{2}(A+2s)\geq 6$, a contradiction.
On the other hand, if $v_{2}(A)\neq 2$, from Equation (7) we obtain
$4=2\min\{v_{2}(A),2\}.$
That is, $2=\min\{v_{2}(A),2\}$, thus $v_{2}(A)\geq 2$. Since $v_{2}(A)\neq 2$, we have $v_{2}(A)\geq 3$. Since $B=s^{2}=4r^{2}$, we have that $v_{2}(B)=1$. Using the expression of $B$ in terms of $\gamma$, we have that
$2=v_{2}(B)=v_{2}(3\gamma^{2}+2\gamma b_{2}+8b_{4})=v_{2}(3\gamma^{2}+2\gamma b_{2})=v_{2}(\gamma)+v_{2}(3\gamma+2b_{2}).$
The above equation implies that $v_{2}(\gamma)\neq 0$, and also that $v_{2}(\gamma)<2$. Therefore, $v_{2}(\gamma)=1$, say $\gamma=2k$ with $k$ an odd integer. Moreover, since both terms are greater than or equal to $1$, we conclude that they are exactly equal to $1$; in particular $1=v_{2}(3\gamma+2b_{2})=v_{2}(6k+2b_{2})=1+v_{2}(3k+b_{2})$. This forces $b_{2}$ to be even. Hence $0<v_{2}(b_{2})=v_{2}(a_{1}^{2}+4a_{2})$ (where we have used [11, III.1 p.42] for the formula of $b_{2}$). Recalling that $a_{1}\in\{0,1\}$, this implies that $a_{1}=0$ and we get
$3\leq v_{2}(A)=v_{2}(3\gamma+4a_{2})=1,$
a contradiction.
I.C: $v_{2}(s)=0$. As $B=s^{2}$, we have that
$2\nmid B=3\gamma^{2}+2\gamma b_{2}+8b_{4},$
so $2\nmid\gamma$. Moreover from Equation (6) we obtain $8=v_{2}(A+2s)+v_{2}(A-2s)$, therefore $A$ must be even. But $2|A=3\gamma+b_{2}$, so $b_{2}$ is odd. From [11, III.1 p.42] we know that
$b_{2}$ $=$ $a_{1}^{2}+4a_{2},$
$b_{4}$ $=$ $2a_{4}+a_{1}a_{3},$
$b_{6}$ $=$ $a_{3}^{2}+4a_{6}.$
We have assumed $a_{1},a_{3}\in\{0,1\}$ and $a_{2}\in\{-1,0,1\}$. If $a_{1}=0$, $b_{2}$ is even, a contradiction. Therefore, $a_{1}=1$. We also have the equation
$8=v_{2}(A-2s)+v_{2}(A+2s).$
Note that both the terms of the right-hand side must be odd. First, $2|d$ and $d$ is squarefree. And, of course, any other integer $d^{\prime}$ such that $\mathbb{Q}(\sqrt{d})=\mathbb{Q}(\sqrt{d^{\prime}})$ must satisfy $d/d^{\prime}\in(\mathbb{Q}^{*})^{2}$. So, as $\mathbb{Q}(\sqrt{d})=\mathbb{Q}(\sqrt{A\pm 2s})$, both $v_{2}(A-2s)$ and $v_{2}(A+2s)$ must be odd.
Additionally, we know
$\gamma^{3}+b_{2}\gamma^{2}+8b_{4}\gamma+16b_{6}=0,$
so $\gamma^{2}(\gamma+b_{2})=-8(b_{4}\gamma+2b_{6})$. Therefore, $v_{2}(\gamma+b_{2})\geq 3$ and
$v_{2}(A)=v_{2}(2\gamma+(\gamma+b_{2}))=1.$
Say $A=2M$ with $M$ odd. Then $v_{2}(A\pm 2s)=v_{2}(2(M\pm s))\geq 2$. As $v_{2}(A\pm 2s)$ is an odd integer, we have that $v_{2}(A\pm 2s)\geq 3$. In this way,
$3\leq v_{2}(A\pm 2s)=v_{2}(2(\gamma\pm s)+(\gamma+b_{2})).$
Hence $v_{2}(\gamma\pm s)\geq 2$, but
$1=v_{2}(2s)=v_{2}((s+\gamma)+(s-\gamma))\geq 2,$
a contradiction.
So with the hypothesis of this subsection, if $E(\mathbb{Q})[2]=\mathcal{C}_{2}$ and $2|d$, the curve $E$ has bad reduction at 2.
Case II: $E_{\mathrm{tors}}(\mathbb{Q})[2]=\mathcal{C}_{2}\times\mathcal{C}_{2}$. For this purpose, we will use the following classical result in the literature of elliptic curves [7, Thm. 4.2].
@@
with $\alpha,\beta,\gamma\in L$. For $P=(x_{0},y_{0})\in E(L)$, there exists $Q\in E(L)$ such that $[2]Q=P$ if and only if $x_{0}-\alpha,x_{0}-\beta,x_{0}-\gamma\in\left(L^{*}\right)^{2}$.
@@
with $\alpha,\beta\in\mathbb{Q}$ and we know that there exists $Q\in E(\mathbb{Q}(\sqrt{d}))\setminus E(\mathbb{Q})$ such that $[2]Q=(0,0)$. Thus
$-\alpha,-\beta\in\left(K^{*}\right)^{2}.$
Note that $-\alpha$ and $-\beta$ cannot be both squares in $\mathbb{Q}$, because in that case
@@
Besides, there exists a point $Q\in E(\mathbb{Q})[4]$ and the four points of $E[4]$ that map to $(0,0)$ under multiplication by $2$ are precisely $Q,Q+(0,0),Q+P_{1},Q+P_{2}$. These are all rational, which contradicts our hypothesis that there is some point $P$ of $4$-torsion which is not rational and such that $[2]P=(0,0)$.
@@
II.A: $-\alpha=a^{2}d,-\beta=b^{2}$. Equation (6) becomes
@@
11 11 $=$ $v_{2}(3\gamma+b_{2})$
33 33 $=$ $v_{2}(3\gamma^{2}+2\gamma b_{2}+8b_{4}).$
@@
So $v_{2}(b)=0$ and $01=v_{2}(a^{2}d-b^{2})=$ 0, a contradiction.
II.B: $-\alpha=a^{2},-\beta=b^{2}d$. This case is analogous to the previous one.
II.C: $-\alpha=a^{2}d,-\beta=b^{2}d$. Now we have the following equations:
$A$ $=$ $d(a^{2}+b^{2})$
$B$ $=$ $a^{2}b^{2}d^{2}$
$2^{12}\Delta$ $=$ $16a^{4}b^{4}d^{6}(a^{2}-b^{2})^{2}$
@@
It is a natural question if this statement still holds when we change the hypothesis of $2|d$ by $2$ ramifies in $K$. The answer is that it does not hold, a counterexample is the curve $217.a$ 2.
@@
Now we address the case where we have an elliptic curve $E/\mathbb{Q}$ with a point $P\in E(K)\setminus E(\mathbb{Q})$ of order $8$ or $16$ for a quadratic extension $K=\mathbb{Q}(\sqrt{d})$, such that $[2]P\in E(\mathbb{Q})$.
@@
where we have the following relation between discriminants
@@
and where, by the Nagell-Lutz theorem, the point $Q:=[2]P=(x_{1},y_{1})\in E_{2}(\mathbb{Q})$ has integer coordinates.
A Tate normal form is an equation like
@@
$E_{3}:y^{2}+\overline{a}_{1}xy+\overline{a}_{3}y=x^{3}+\overline{a}_{2}x^{2}$
Every change of variables that preserves the Weistrass form must be like
$x=u^{2}x^{\prime}+r,\quad y=u^{3}y^{\prime}+su^{2}x^{\prime}+t$
@@
we require that $r=x_{1}$ and $t=y_{1}$. This makes 00 the independent term of the equation. Now, we choose $s$ in order to force the coefficient in $x$ to be 00, which give us the following relation
@@
$\left\{\begin{array}[]{lcl}\overline{a}_{1}&=&2s\\ \overline{a}_{2}&=&-s^{2}+3x_{1}+b_{2}\\ \overline{a}_{3}&=&2y_{1}\\ \Delta_{3}&=&2^{12}\Delta_{1}\end{array}\right.$
Note that $\overline{a}_{2}=0$ if and only if the point $Q$ is a $3$-torsion point (see [7, V.5]). In order to get a Tate normal form, we just have to equalize the coefficients of $y$ and $x^{2}$. We get the equation
@@
If we call $b:=-\tilde{a}_{2}=-\tilde{a}_{3}$ and $c:=1-\tilde{a}_{1}$, we get the usual Tate normal form
@@
Let $E/\mathbb{Q}$ be an elliptic curve and $K=\mathbb{Q}(\sqrt{d})$ with $d$ squarefree and such that $2|d$. If there exists $P\in E(K)[8]$ of order 8 such that $[2]P\in E(\mathbb{Q})[4]$, then $E$ has bad reduction at $2$.
@@
We go as usual by contradiction, assuming that $E$ has good reduction at $2$, i.e. $\Delta_{1}$ is an odd integer. In order to get a Tate normal form, we use the previous coordinates changes. Since $Q=[2]P$ has order $4$, $c$ must be 00 (see [4, Chapter 4, §$\S$ 4]). So our Tate normal form is
(11) $\mathcal{T}_{b,0}:y^{2}+xy-by=x^{3}-bx^{2}.$
@@
    $E_{\mathrm{tors}}(\mathbb{Q})={\mathcal{C}}_{8}$. The only growths that can occur (see Theorem $2$ of [2]) are $E_{\mathrm{tors}}(K)={\mathcal{C}}_{16}$ or $E_{\mathrm{tors}}(K)={\mathcal{C}}_{2}\times{\mathcal{C}}_{8}$. In the first case, there are no new points of order $8$, but in the second case we obtain points of order $8$ which are defined only after a base change to $K$. This case is considered in section $4.1$ of [3]; the only possibility for $K$ is $K=\mathbb{Q}(\sqrt{\Delta_{E}})$, which contradicts the condition $2|d$.
@@
    $E_{\mathrm{tors}}(\mathbb{Q})={\mathcal{C}}_{12}$. According to theorem $2$ of [2], the only possible growth is to ${\mathcal{C}}_{2}\times{\mathcal{C}}_{12}$. The same reasoning as above shows that this cannot happen.
@@
Thus, we have two possibilities: either $E_{\mathrm{tors}}(\mathbb{Q})={\mathcal{C}}_{4}$ or $E_{\mathrm{tors}}(\mathbb{Q})={\mathcal{C}}_{2}\times{\mathcal{C}}_{4}$. In both cases, we are going to apply Lemma $14$ of [3].
@@
Let $E$ be an elliptic curve defined over $\mathbb{Q}$ with $E(\mathbb{Q})_{\mathrm{tors}}=\mathcal{C}_{4}$. Let $t\in\mathbb{Q}$ such that $E$ is $\mathbb{Q}$–isomorphic to $\mathcal{T}_{t,0}$. There exists a quadratic field $K$ with $E(K)_{\mathrm{tors}}=\mathcal{C}_{8}$ if and only if $t=-s^{2}$ for some $s\in\mathbb{Q}$.
@@
In this result the hypothesis states $E_{\mathrm{tors}}(\mathbb{Q})={\mathcal{C}}_{4}$. However, in Section $4.2$ of [3], the authors explain that the first part of the proof of this lemma is valid also in the case when $E_{\mathrm{tors}}(\mathbb{Q})={\mathcal{C}}_{2}\times{\mathcal{C}}_{4}$ which is the part we actually need.
Then we have that there exists $r\in\mathbb{Q}$ such that $b=-r^{2}$ and $K=\mathbb{Q}(\sqrt{1\pm 4r})$. Let us write $r=p/q$ with $\gcd(p,q)=1$. So $\mathbb{Q}(\sqrt{d})=\mathbb{Q}(\sqrt{q(q\pm 4p)})$, which implies that $2|q(q\pm 4p)$. Hence $2|q$ and $2\nmid p$. We will look closely first at the precise value of $v_{2}(q)$.
@@
and implies that $3$ is even, a contradiction.
@@
(13) $-\frac{p^{2}}{q^{2}}=b=-\frac{\overline{a}_{2}^{3}}{\overline{a}_{3}^{2}}=-\frac{\overline{a}_{3}}{\overline{a}_{1}^{3}},$
where we have used that $c=0$ in the last equality. Taking the $2$-adic valuation and using the definition of $\overline{a}_{i}$, we get
@@
Now let us assume that $v_{2}(q)\geq 3$. In this way the last equation becomes
@@
which implies that $v_{2}(q)$ is even. However there exists odd numbers $M,N$ such that $q=2^{v_{2}(q)}N$ and $q\pm 4p=4M$. But this implies that
@@
Therefore we can affirm $v_{2}(q)=2$ and the previous equation becomes
@@
where the last inequality comes from the fact that $v_{2}(q\pm 4p)\geq 3$, as we know that $v_{2}(p)=0$ and $v_{2}(q)=2$. The inequality implies that $v_{2}(y_{1})\in\{0,1\}$. If $v_{2}(y_{1})=0$, we have
$-4=v_{2}(b)=3v_{2}(\overline{a}_{2})-2v_{2}(\overline{a}_{3})=3v_{2}(\overline{a}_{2})-2.$
[…]
$0=v_{2}(1)=v_{2}(\overline{a}_{1})+v_{2}(\overline{a}_{2})-v_{2}(\overline{a}_{3}).$

### v2

@@ 3. The prime $\ell=2$: The mixed case
Let us continue with the case where a new torsion point $P\in E(K)$ appears, such that a nontrivial multiple $mP$ belongs to $E(\mathbb{Q})$. As mentioned in the Introduction, from Theorem 1, we can reduce ourselves to the following hypothesis:
$\text{There exists $P\in E(K)[N]\setminus E(\mathbb{Q})$ such that $2P\in E(\mathbb{Q})$ and $o(P)=N$},$
@@
Let $E/\mathbb{Q}$ be an elliptic curve and $K=\mathbb{Q}(\sqrt{d})$ a quadratic extension. Let us suppose that there exists $P\in E(K)[4]\setminus E(\mathbb{Q})$ such that $\mathcal{O}\neq 2P\in E(\mathbb{Q})$. Then if $2|d$, $E$ has bad reduction at $2$.
@@
where formulas for $b_{i}$ can be found in [14, Chapter III]. It is easy to see that the discriminant of this equation is still $\Delta$. In order to get an equation with integer coefficients, we make the change
@@
By hypothesis, $\mathcal{O}\neq 2P\in E_{\mathrm{tors}}(\mathbb{Q})$. So there exists $2P=Q=(\gamma,0)$ with $\gamma\in\mathbb{Z}$ by Nagell-Lutz theorem. Now, we apply the change of variables
@@
Case I: $E_{\mathrm{tors}}(\mathbb{Q})[2]=\mathcal{C}_{2}$. We can use the following lemma proven in [4, Lemma 13].
@@
At least one among $v_{2}(A-2s),v_{2}(A+2s)$ is odd as $v_{2}(d)$ is odd. Hence, they must be both odd and distinct, as if they were the same, their sum would not be divisible by 4. For $v_{2}(A-2s)$ and $v_{2}(A+2s)$ to be distinct, we must have $v_{2}(A)=v_{2}(2s)$. Furthermore, $v_{2}(A-2s)$ and $v_{2}(A+2s)$ are two odd integer greater than $v_{2}(2s)$, so their sum is at least $2v_{2}(s)+4=2v_{2}(s)+6$, which implies $v_{2}(s)=0$, and furthermore $v_{2}(A)=1$.
Writing $A=2r$ with $r$ odd, we have
$8=2+v_{2}(r-s)+v_{2}(r+s),$
with $v_{2}(r-s)$ and $v_{2}(r+s)$ even and positive. However, we cannot have both rsr-s and $r+s$ divisible by $4$, since $r$ and $s$ are odd. Thus, at least one among $v_{2}(r-s),v_{2}(r+s)$ equals $1$, a contradiction.
Case II: $E_{\mathrm{tors}}(\mathbb{Q})[2]=\mathcal{C}_{2}\times\mathcal{C}_{2}$. For this purpose, we will use the following classical result in the literature of elliptic curves [8, Theorem 4.2].
@@
with $\alpha,\beta,\gamma\in L$. For $P=(x_{0},y_{0})\in E(L)$, there exists $Q\in E(L)$ such that $2Q=P$ if and only if $x_{0}-\alpha,x_{0}-\beta,x_{0}-\gamma\in\left(L^{*}\right)^{2}$.
@@
with $\alpha,\beta\in\mathbb{Q}$ and we know that there exists $Q\in E(\mathbb{Q}(\sqrt{d}))\setminus E(\mathbb{Q})$ such that $2Q=(0,0)$. Thus
$\alpha,-\beta\in\left(\mathbb{Q}(\sqrt{d})^{*}\right)^{2}.$
Note that $\alpha$ and $\beta$ cannot be both squares in $\mathbb{Q}$. Should that be the case, let us write
@@
Then from the previous lemma, as $\alpha$ and $\beta$ a rational squares, there exists a point $R\in E(\mathbb{Q})[4]$ and the four points of $E[4]$ that map to $(0,0)$ under multiplication by $2$ are precisely $R,R+(0,0),R+P_{1},R+P_{2}$. These are all rational, which contradicts our hypothesis that there is some point $Q$ of $4$-torsion which is not rational and such that $2Q=(0,0)$.
@@
II.A: $\alpha=a^{2}d,-\beta=b^{2}$. Equation (6) becomes
@@
$1$ $=$ $v_{2}(3\gamma+b_{2}),$
$3$ $=$ $v_{2}(3\gamma^{2}+2\gamma b_{2}+8b_{4}).$
@@
So $v_{2}(b)=0$ and $1=v_{2}(a^{2}d-b^{2})=0$, a contradiction.
II.B: $\alpha=a^{2},-\beta=b^{2}d$. This case is analogous to the previous one.
II.C: $\alpha=a^{2}d,-\beta=b^{2}d$. Now we have the following equations:
$A$ $=$ $d(a^{2}+b^{2}),$
$B$ $=$ $a^{2}b^{2}d^{2},$
$2^{12}\Delta$ $=$ $16a^{4}b^{4}d^{6}(a^{2}-b^{2})^{2}.$
@@
Remark 2.
It is a natural question if this statement still holds when we change the hypothesis of $2|d$ by the weaker hypothesis that $2$ ramifies in $K$. The answer is that it does not hold, a counterexample is the curve $17.a2$.
@@
Now we address the case where we have an elliptic curve $E/\mathbb{Q}$ with a point $P\in E(K)\setminus E(\mathbb{Q})$ of order $8$ or $16$ for a quadratic extension $K=\mathbb{Q}(\sqrt{d})$, such that $2P\in E(\mathbb{Q})$.
@@
where we have the following relation between the respective discriminants of $E_{1}$ and $E_{2}$
@@
and where, by the Nagell-Lutz theorem, the point $Q=2P=(x_{1},y_{1})\in E_{2}(\mathbb{Q})$ has integer coordinates.
A Tate normal form is an equation of the form
@@
$E_{3}:y^{2}+\overline{a}_{1}xy+\overline{a}_{3}y=x^{3}+\overline{a}_{2}x^{2}.$
Every change of variables that preserves the Weierstrass form must be like
$x=u^{2}x^{\prime}+r,\quad y=u^{3}y^{\prime}+su^{2}x^{\prime}+t.$
@@
we require that $r=x_{1}$ and $t=y_{1}$. This makes 0 the independent term of the equation. Now, we choose $s$ in order to force the coefficient in $x$ to be 0, which give us the following relation
@@
$\left\{\begin{array}[]{lcl}\overline{a}_{1}&=&2s,\\ \overline{a}_{2}&=&-s^{2}+3x_{1}+b_{2},\\ \overline{a}_{3}&=&2y_{1},\\ \Delta_{3}&=&2^{12}\Delta_{1}.\end{array}\right$.
Note that $\overline{a}_{2}=0$ if and only if the point $Q$ is a $3$-torsion point (see [8, V.5]). In order to get a Tate normal form, we just have to equalize the coefficients of $y$ and $x^{2}$. We get the equation
@@
If we call $b=-\tilde{a}_{2}=-\tilde{a}_{3}$ and $c=1-\tilde{a}_{1}$, we get the usual Tate normal form
@@
Let $E/\mathbb{Q}$ be an elliptic curve and $K=\mathbb{Q}(\sqrt{d})$ with $d$ a squarefree even integer. If there exists a point $P\in E(K)[8]$ of order 8 such that $2P\in E(\mathbb{Q})$ then $E$ has bad reduction at $2$.
@@
We go as usual by contradiction, assuming that $E$ has good reduction at $2$, i.e. $\Delta_{1}$ is an odd integer. In order to get a Tate normal form, we use the previous coordinates changes. Since $Q=2P$ has order $4$, $c$ must be 0 (see [5, Chapter 4, $\S$4]). So our Tate normal form is
(10) $\mathcal{T}_{b,0}:y^{2}+xy-by=x^{3}-bx^{2},$
and, furthermore, we must have $1=\tilde{a}_{1}$ and hence $\overline{a}_{1}\overline{a}_{2}=\overline{a}_{3}$.
@@
    $E_{\mathrm{tors}}(\mathbb{Q})={\mathcal{C}}_{8}$. The only growths that can occur (see Theorem 1) are $E_{\mathrm{tors}}(K)={\mathcal{C}}_{16}$ or $E_{\mathrm{tors}}(K)={\mathcal{C}}_{2}\times{\mathcal{C}}_{8}$. In the first case, there are no new points of order $8$, but in the second case we obtain points of order $8$ which are defined only after a base change to $K$. This case is considered in [4, Section 4.1]; the only possibility for $K$ is $K=\mathbb{Q}(\sqrt{\Delta_{E}})$, which contradicts the condition $2|d$.
@@
    $E_{\mathrm{tors}}(\mathbb{Q})={\mathcal{C}}_{12}$. According to Theorem 1, the only possible growth is to ${\mathcal{C}}_{2}\times{\mathcal{C}}_{12}$. The same reasoning as above shows that this cannot happen.
@@
Thus, we have two possibilities: either $E_{\mathrm{tors}}(\mathbb{Q})={\mathcal{C}}_{4}$ or $E_{\mathrm{tors}}(\mathbb{Q})={\mathcal{C}}_{2}\times{\mathcal{C}}_{4}$. In both cases, we are going to apply [4, Lemma 14].
@@
Let $E$ be an elliptic curve defined over $\mathbb{Q}$ with $E_{\mathrm{tors}}(\mathbb{Q})=\mathcal{C}_{4}$. Let $t\in\mathbb{Q}$ such that $E$ is $\mathbb{Q}$–isomorphic to $\mathcal{T}_{t,0}$. There exists a quadratic field $K$ with $E_{\mathrm{tors}}(K)=\mathcal{C}_{8}$ if and only if $t=-s^{2}$ for some $s\in\mathbb{Q}$.
@@
In this result the hypothesis states $E_{\mathrm{tors}}(\mathbb{Q})={\mathcal{C}}_{4}$. However, in [4, Section 4.2], the authors explain that the first part of the proof of this lemma is valid also in the case when $E_{\mathrm{tors}}(\mathbb{Q})={\mathcal{C}}_{2}\times{\mathcal{C}}_{4}$ which is the part we actually need.
Then we have that there exist $r\in\mathbb{Q}$ such that $b=-r^{2}$ and $K=\mathbb{Q}(\sqrt{1\pm 4r})$. Let us write $r=p/q$ with $\gcd(p,q)=1$. So $\mathbb{Q}(\sqrt{d})=\mathbb{Q}(\sqrt{q(q\pm 4p)})$, which implies that $2|q(q\pm 4p)$. Hence $2|q$ and $2\nmid p$. We will look closely first at the precise value of $v_{2}(q)$.
@@
which leads to $5=3\left(v_{2}(8b_{4}+2b_{2}x_{1}+3x_{1}^{2})-v_{2}(y_{1})\right)$, a contradiction.
@@
(12) $\frac{p^{2}}{q^{2}}=b=-\frac{\overline{a}_{2}^{3}}{\overline{a}_{3}^{2}}=-\frac{\overline{a}_{3}}{\overline{a}_{1}^{3}},$
as $\overline{a}_{1}\overline{a}_{2}=\overline{a}_{3}$ in this case. Taking the $2$-adic valuation and using the definition of $\overline{a}_{i}$, we get
@@
Now let us assume that $v_{2}(q)\geq 3$, so $2v_{2}(q)+8=4v_{2}(y_{1})+4$. Dividing by $2$, the equation (14) becomes
@@
which implies that $v_{2}(q)$ is even. However there exist odd numbers $M,N$ such that $q=2^{v_{2}(q)}N$ and $q\pm 4p=4M$. But this implies that
@@
Therefore we can affirm $v_{2}(q)=2$. Substituting it in the equation (13), it becomes
@@
where the last inequality comes from the fact that $v_{2}(q\pm 4p)\geq 3$, as we know that $v_{2}(p)=0$ and $v_{2}(q)=2$. The inequality implies that $v_{2}(y_{1})\in\{0,1\}$, so we can separate the proof in two cases.
Case I: $v_{2}(y_{1})=0$. If $v_{2}(y_{1})=0$, taking the $2$-adic valuation in the equation (12), we get
$4=v_{2}(b)=3v_{2}(\overline{a}_{2})-2v_{2}(\overline{a}_{3})=3v_{2}(\overline{a}_{2})-2.$
Hence, $2=3v_{2}(\overline{a}_{2})$, a contradiction.
Case II: $v_{2}(y_{1})=1$. In this case, $v_{2}(\overline{a}_{3})=1+v_{2}(y_{1})=2$. Now, from Equation (12), we obtain that $2v_{2}(q)=3v_{2}(\overline{a}_{2})-2v_{2}(\overline{a}_{3})$. But in our case $v_{2}(q)=2$, so we must have $v_{2}(\overline{a}_{2})=0$. As $\overline{a}_{1}\overline{a}_{2}=\overline{a}_{3}$,
$0=v_{2}(\overline{a}_{1})+v_{2}(\overline{a}_{2})-v_{2}(\overline{a}_{3}).$